About how many times more time does a brute force key search take against a 112-bit DES than against a 56-bit DES?
Answer #1

So we calculate the key space sizes , that is for 112bit DES = 2^112 , for 56 bit DES is 2^56.
Now we have to find the ratio 2^112/2^56 = 2^56 ≈ 7·10^16 , this times more time it takes for the given here 112-bit DES than 56-bit DES when we go for the brute force.